# FATE: A Formal Benchmark Series for Frontier Algebra of Multiple Difficulty Levels

### v2

@@ Acknowledgments
We wish to express our sincere appreciation to Yutong Wang for generously providing the code for the Lean verification. We are deeply grateful to the following individuals for their contributions to the selection of mathematical problems for the FATE benchmarks: Kaiyi Chen, Haocheng Fan, Yiqin He, Yongle Hu, Shanxiao Huang, Yudong Liu, Tian Qiu, Yinchong Song, Peihang Wu, Zhenhua Wu, Tianyi Xu, Zhehan Xu, Huanhuan Yu, Huishi Yu, Jiahong Yu, Zhanhao Yu and Xiao Yuan. We gratefully acknowledge the following individuals for their careful assessment of natural language proofs and insightful feedback: Haocheng Fan, Haoda Li, Zhongjin Yan, Chaodong Yang, Yuji Yang, Shun Yin, Huanhuan Yu, Jiahong Yu and Weijun Yuan. We also thank Nailin Guan, Zixun Guo, Yongle Hu, and Jingting Wang for their dedicated work in formalizing mathematical problems. We extend our sincere thanks to Johan Commelin and Christian Merten for their careful review and corrections to the benchmark problems. We are also indebted to Siddhartha Gadgil, Filippo A. E. Nuccio and Floris van Doorn for their valuable discussions and suggestions.
@@ Appendix A Introduction to Formalization and Lean
Various ITPs are built on distinct but equally sound logical foundations: Metamath (Megill & Wheeler, 2019) and Mizar (Bancerek et al., 2015; Grabowski et al., 2010) are based on set theory; HOL4 (Slind & Norrish, 2008), HOL Light (Harrison, 2009), and Isabelle/HOL (Nipkow et al., 2002) are based on simple type theory; and Coq (Barras et al., 1999) and Lean (Moura & Ullrich, 2021b) are based on dependent type theory. The mathematical library available in each prover determines the difficulty of formalizing different topics. Lean’s extensive and unified Mathlib (mathlib Community, 2020) distinguishes it from other ITPs.
@@
Lean 4 (Moura & Ullrich, 2021b) is an interactive theorem prover based on dependent type theory, with proof irrelevance and non-cumulative universes (Carneiro, 2019). An introduction to the language can be found in (Avigad et al., 2021). Lean 4 is designed to be highly extensible through its metaprogramming capabilities, enabling the accurate and concise extraction of important metadata.
@@ Appendix B Benchmark Quality Analysis
This appendix presents the quality of the FATE benchmarks, emphasizing their mathematical content. Following illustrative examples from each benchmark in Section˜B.1, we provide case studies and statistics to demonstrate the diversity (see Section˜B.2) and progressive difficulty (see Section˜B.3) of the problems. The results of an expert survey assessing the benchmarks are reported in Section˜B.4. For comparison, examples from the FATE-M benchmark introduced by Shen et al. (2025) are included. For analyses focused purely on mathematical content, we provide only the natural-language problem statements without formalization details for brevity.
@@ B.1 Representative Examples
Suppose \$D\$ is integral domain, \$m\$ and \$n\$ are coprime positive integers.
@@
Let $\mathbb${F}_4\$ be the field with \$4\$ elements, \$t\$ a transcendental over
$\mathbb${F}_4\$, and \$F $\mathbb${F}_4(t^4 + t)\$ and \$K $\mathbb${F}_4(t)\$.
Show that \$K\$ is Galois over \$F\$.
@@
Let \( A \) be a domain and \( K \) its field of fractions. \( x $\in$ K \) is
@@
\( A \) is called completely integrally closed if every almost integral element
of \( K \) is contained in \( A \).
def IsCompletelyIntegrallyClosed (A : Type) [CommRing A] [IsDomain A] : Prop :=
x : FractionRing A, IsAlmostIntegral x
y : A, x = algebraMap A (FractionRing A) y
Let \( A \) be a domain. Show that if \( A \) is completely integrally closed,
so is \( A[X] \).
@@ B.2 Diversity
Section˜B.2.1 presents three principal problem types within the series, each accompanied by a representative case. The statistical distribution of mathematical domains is detailed in Section˜B.2.2.
@@ B.2.1 Problem Types
The first category, abstract reasoning problems, is the most common type in FATE-H and FATE-X. It is noteworthy that the abstractness of a problem type does not directly correlate with its mathematical difficulty. A detailed case study on the benchmark’s difficulty gradient is available in Section˜B.3.1.
@@ B.2.2 Domain Statistics
We classify abstract algebra and commutative algebra into subfields and further break down these subfields into specific topics. For a complete list of subfields, topics, and corresponding statistics, see Table˜4; for a visual sunburst chart representation, refer to Figure˜3. The chart illustrates a shift in emphasis from abstract algebra in FATE-H to advanced commutative algebra topics in FATE-X.
@@
Dimension Theory
and Smoothness
Heights and Dimensions 2 6
@@ B.3 Progressive Difficulty
Section˜B.3.1 presents three topic-related examples of increasing difficulty, demonstrating the progressive structure of the benchmarks. Furthermore, Section˜B.3.2 provides statistics on model output token length and natural language accuracy to reflect problem difficulty.
@@ B.3.1 Case Study: A Graded Challenge on Non-Simple Groups
To demonstrate the difficulty gradient of the FATE benchmark, we analyze the solutions and reasoning contexts generated by DeepSeek-R1 on three topic-related problems from FATE-M, H, and X. As illustrated in Figure˜4, the model’s generated proofs exhibit a clear progression in structure and complexity, reflecting the increasing mathematical difficulty of the problems.
@@
The increasing complexity of the required proof structure is directly reflected in the model’s output. The reasoning content produced by the model aligns with this structural progression, as visualized in Figure˜4 using distinct colors to mark different reasoning stages:
@@ B.3.2 Auxiliary Statistics on Proof Difficulty
As noted at the end of Section˜4.2, the model first produces a complete natural language solution before formalizing it. The data in Table˜5 reflect expert evaluations of these intermediate natural language solutions (Pass@1), showing a clear decreasing accuracy trend from FATE-M to FATE-H to FATE-X.
@@
It should be noted that these results do not directly represent natural language performance, since the model’s final task is formalization rather than natural language proof generation. An ablation study in Section˜C.2 compares natural language accuracy between a direct natural language task and the intermediate natural language output within the formalization task. The results show that although a pure natural language task yields higher accuracy, the improvement is significantly smaller than the gap between different levels and therefore does not affect the observed decreasing accuracy trend across the three benchmarks.
@@
The token length data in Table˜6 indicate a clear increasing trend in output token length from FATE-M to FATE-H to FATE-X across all models.
@@ B.4 Expert Assessment
In the first part, we divided the FATE-H and FATE-X benchmarks, classical algebra textbooks Dummit et al. (2004), and the algebra component of ProofNet by mathematical topic. We then randomly sampled 10 problems with balanced topic distribution from each to form three anonymous problem sets. Participants were asked to compare and select the best set across several dimensions. The results are summarized as follows:
@@ C.2 Ablation Study of Prompts
To understand how the prompt influence the model’s natural language performance, we designed a pass@1 ablation study testing four different prompt settings for DeepSeek-R1 on FATE-X, see Section˜G.2 for explicit prompts:
@@
The results, shown in Table˜7, reveal the following:
@@ C.3.4 Reasoning Problem
Unlike the previous three categories, which relate primarily to mathematical knowledge or strategy, a Reasoning Problem stems from deficiencies in the model’s general-purpose reasoning capabilities. Here, the model uses largely correct mathematical facts but connects them with a flawed logical flow. Because the patterns of these errors are more complex, we provide a detailed discussion and case analysis in Appendix˜E.
@@ Appendix D Analysis of Formal Language Output
This appendix provides an in-depth characterization of the four formalization error categories identified in Section˜4.4, accompanied by illustrative case studies.
@@ D.1 Detailed Formal Errors
have h : Fintype.card (Sylow 3 G) 1 [MOD 3] := by
have h : Fintype.card (Sylow 3 G) 1 [MOD 3] := by
have h : Fintype.card (Sylow 3 G) 1 [MOD 3] := by
have h ₁₀ : Fact (Nat.Prime 3) := h
have h ₁₁ : Fintype.card (Sylow 3 G) 1 [MOD 3] := by
@@
exact h ₁₁
exact h
exact h
@@
have h : IsIntegral (sin / 180)) := by
have h : IsIntegral (sin / 180)) := by
@@
have h : IsAlgebraic (sin / 180)) := h₁.isAlgebraic
@@ D.2 Case study on FATE-X
A comparison of intermediate natural language accuracy on FATE-H (Table˜8) reveals that DeepSeek-R1 significantly outperforms both DeepSeek-V3 and DeepSeek-Prover-V2. The latter two models exhibit similar performance.
@@ E.1 Analysis of Behaviors during the Reasoning Process
Therefore, the ideal generated by X and Y <...> No, because 1 = - (X2+Y2) <...> For example, the point (i,0) is defined over the complexes <...> we cannot write it as (X - i, Y) because i is not in the base ring. <...> So (X, Y-1) has no common zero, so by Hilberts Nullstellensatz, it is the unit ideal?
@@
Analysis: This solution is also discussed in the most difficult example of Section˜B.3.1. As the analysis before, after applying sylow theorem and counting elements, the model reduces the problem to eight possible combinations of Sylow subgroup counts. It then explores a specific case (Combination 1) and derives an intermediate result—that the normalizer of a Sylow 3-subgroup has order 255—but this path does not lead to a contradiction. The model then returns to the general case and successfully deduces a key relationship between the Sylow 17-subgroups and Sylow 3-subgroups. Unfortunately, it then incorrectly applies the intermediate result (order 255), which is valid only for Combination 1, to the general argument, causing the proof to fail.
@@ Appendix F Experiment Details
In Section˜F.1, we detailed our formal experiment setup, encompassing both the generation and verification phases, while Section˜F.2 further elaborates on the process, standards, and methodology for our manual natural language evaluations.
@@ F.1.2 Verification
This appendix explained the verification methodology in Section˜4.1 employed to ensure the correctness and integrity of the generated Lean proofs, especially considering the models’ flexibility to generate auxiliary definitions and lemmas.
@@
Upon successfully passing all the string-based checks, every formal proof is rigorously checked by the Lean kernel to confirm it contains no sorry or compilation errors.
@@ F.1.3 Additional Pass@k Results
We report Pass@$k$ results (with $k=2^{n}$ for $n=0,\dots,6$) in Table˜9 for the FATE-M and FATE-H benchmarks, omitting FATE-X as no model achieved a successful proof. These results are estimated from $n=64$ samples per problem using the unbiased estimator:
@@
Table 9: Pass@k results across the FATE series
(a) Pass@k on FATE-M (%)

### v3

@@ Acknowledgments
We wish to express our sincere appreciation to Yutong Wang for generously providing the code for the Lean verification. We are deeply grateful to the following individuals for their contributions to the selection of mathematical problems for the FATE benchmarks: Kaiyi Chen, Haocheng Fan, Yiqin He, Yongle Hu, Shanxiao Huang, Yudong Liu, Tian Qiu, Yinchong Song, Peihang Wu, Zhenhua Wu, Tianyi Xu, Zhehan Xu, Huanhuan Yu, Huishi Yu, Jiahong Yu, Zhanhao Yu and Xiao Yuan. We gratefully acknowledge the following individuals for their careful assessment of natural language proofs and insightful feedback: Haocheng Fan, Haoda Li, Zhongjin Yan, Chaodong Yang, Yuji Yang, Shun Yin, Huanhuan Yu, Jiahong Yu and Weijun Yuan. We also thank Nailin Guan, Zixun Guo, Yongle Hu, and Jingting Wang for their dedicated work in formalizing mathematical problems. We extend our sincere thanks to Johan Commelin and Christian Merten for their careful external review and corrections to the benchmark problems. We are also indebted to Siddhartha Gadgil, Filippo A. E. Nuccio and Floris van Doorn for their valuable discussions and suggestions.
@@ Appendix A Introduction to Formalization and Lean
Various ITPs are built on distinct but equally sound logical foundations: Metamath (Megill & Wheeler, 2019) and Mizar (Bancerek et al., 2015; Grabowski et al., 2010) are based on set theory; HOL4 (Slind & Norrish, 2008), HOL Light (Harrison, 2009), and Isabelle/HOL (Nipkow et al., 2002) are based on simple type theory; and Coq (Barras et al., 1999) and Lean (Moura & Ullrich, 2021) are based on dependent type theory. The mathematical library available in each prover determines the difficulty of formalizing different topics. Lean’s extensive and unified Mathlib (mathlib Community, 2020) distinguishes it from other ITPs.
@@
Lean 4 (Moura & Ullrich, 2021) is an interactive theorem prover based on dependent type theory, with proof irrelevance and non-cumulative universes (Carneiro, 2019). An introduction to the language can be found in (Avigad et al., 2021). Lean 4 is designed to be highly extensible through its metaprogramming capabilities, enabling the accurate and concise extraction of important metadata.
@@ Appendix B Benchmark Quality Analysis
This appendix presents the quality of the FATE benchmarks, emphasizing their mathematical content. Following illustrative examples from each benchmark in Section B.1, we provide case studies and statistics to demonstrate the diversity (see Section B.2) and progressive difficulty (see Section B.3) of the problems. The results of an expert survey assessing the benchmarks are reported in Section B.4. For comparison, examples from the FATE-M benchmark introduced by Shen et al. (2025) are included. For analyses focused purely on mathematical content, we provide only the natural-language problem statements without formalization details for brevity.
@@ B.1 Representative Examples
Let $\mathbb${F}_4\$ be the field with \$4\$ elements, \$t\$ a transcendental over $\mathbb${F}_4\$, and \$F $\mathbb${F}_4(t^4 + t)\$ and \$K $\mathbb${F}_4(t)\$. Show that \$K\$ is Galois over \$F\$.
@@
Let \( A \) be a domain and \( K \) its field of fractions.
\( x $\in$ K \) is called almost integral if there exists an element
@@
\( A \) is called completely integrally closed if
every almost integral element of \( K \) is contained in \( A \).
def IsCompletelyIntegrallyClosed (A : Type) [CommRing A] [IsDomain A] : Prop := x : FractionRing A, IsAlmostIntegral x y : A, x = algebraMap A (FractionRing A) y
Let \( A \) be a domain. Show that if \( A \) is completely
integrally closed, so is \( A[X] \).
@@ B.2 Diversity
Section B.2.1 presents three principal problem types within the series, each accompanied by a representative case. The statistical distribution of mathematical domains is detailed in Section B.2.2.
@@ B.2.1 Problem Types
The first category, abstract reasoning problems, is the most common type in FATE-H and FATE-X. It is noteworthy that the abstractness of a problem type does not directly correlate with its mathematical difficulty. A detailed case study on the benchmark’s difficulty gradient is available in Section B.3.1.
@@ B.2.2 Domain Statistics
We classify abstract algebra and commutative algebra into subfields and further break down these subfields into specific topics. For a complete list of subfields, topics, and corresponding statistics, see Table 4; for a visual sunburst chart representation, refer to Figure 3. The chart illustrates a shift in emphasis from abstract algebra in FATE-H to advanced commutative algebra topics in FATE-X.
@@
Dimension Theory and Smoothness Heights and Dimensions 2 6
@@ B.3 Progressive Difficulty
Section B.3.1 presents three topic-related examples of increasing difficulty, demonstrating the progressive structure of the benchmarks. Furthermore, Section B.3.2 provides statistics on model output token length and natural language accuracy to reflect problem difficulty.
@@ B.3.1 Case Study: A Graded Challenge on Non-Simple Groups
To demonstrate the difficulty gradient of the FATE benchmark, we analyze the solutions and reasoning contexts generated by DeepSeek-R1 on three topic-related problems from FATE-M, H, and X. As illustrated in Figure 4, the model’s generated proofs exhibit a clear progression in structure and complexity, reflecting the increasing mathematical difficulty of the problems.
@@
The increasing complexity of the required proof structure is directly reflected in the model’s output. The reasoning content produced by the model aligns with this structural progression, as visualized in Figure 4 using distinct colors to mark different reasoning stages:
@@ B.3.2 Auxiliary Statistics on Proof Difficulty
As noted at the end of Section 4.2, the model first produces a complete natural language solution before formalizing it. The data in Table 5 reflect expert evaluations of these intermediate natural language solutions (Pass@1), showing a clear decreasing accuracy trend from FATE-M to FATE-H to FATE-X.
@@
It should be noted that these results do not directly represent natural language performance, since the model’s final task is formalization rather than natural language proof generation. An ablation study in Section C.2 compares natural language accuracy between a direct natural language task and the intermediate natural language output within the formalization task. The results show that although a pure natural language task yields higher accuracy, the improvement is significantly smaller than the gap between different levels and therefore does not affect the observed decreasing accuracy trend across the three benchmarks.
@@
The token length data in Table 6 indicate a clear increasing trend in output token length from FATE-M to FATE-H to FATE-X across all models.
@@ B.4 Expert Assessment
In the first part, we divided the FATE-H and FATE-X benchmarks, classical algebra textbooks (Dummit et al., 2004), and the algebra component of ProofNet by mathematical topic. We then randomly sampled 10 problems with balanced topic distribution from each to form three anonymous problem sets. Participants were asked to compare and select the best set across several dimensions. The results are summarized as follows:
@@ C.2 Ablation Study of Prompts
To understand how the prompt influence the model’s natural language performance, we designed a pass@1 ablation study testing four different prompt settings for DeepSeek-R1 on FATE-X, see Section G.2 for explicit prompts:
@@
The results, shown in Table 7, reveal the following:
@@ C.3.4 Reasoning Problem
Unlike the previous three categories, which relate primarily to mathematical knowledge or strategy, a Reasoning Problem stems from deficiencies in the model’s general-purpose reasoning capabilities. Here, the model uses largely correct mathematical facts but connects them with a flawed logical flow. Because the patterns of these errors are more complex, we provide a detailed discussion and case analysis in Appendix E.
@@ Appendix D Analysis of Formal Language Output
This appendix provides an in-depth characterization of the four formalization error categories identified in Section 4.4, accompanied by illustrative case studies.
@@ D.1 Detailed Formal Errors
have h7 : Fintype.card (Sylow 3 G) 1 [MOD 3] := by
have h8 : Fintype.card (Sylow 3 G) 1 [MOD 3] := by
have h9 : Fintype.card (Sylow 3 G) 1 [MOD 3] := by
have h10 : Fact (Nat.Prime 3) := h3
have h11 : Fintype.card (Sylow 3 G) 1 [MOD 3] := by
@@
exact h11
exact h9
exact h8
@@
have h1 : IsIntegral (sin / 180)) := by
have h2 : IsIntegral (sin / 180)) := by
@@
have h2 : IsAlgebraic (sin / 180)) := h1.isAlgebraic
@@ D.2 Case study on FATE-X
A comparison of intermediate natural language accuracy on FATE-H (Table 8) reveals that DeepSeek-R1 significantly outperforms both DeepSeek-V3 and DeepSeek-Prover-V2. The latter two models exhibit similar performance.
@@ E.1 Analysis of Behaviors during the Reasoning Process
Therefore, the ideal generated by X and Y <...> No, because 1 = - (X²+Y²) <...> For example, the point (i,0) is defined over the complexes <...> we cannot write it as (X - i, Y) because i is not in the base ring. <...> So (X, Y-1) has no common zero, so by Hilberts Nullstellensatz, it is the unit ideal?
@@
Analysis: This solution is also discussed in the most difficult example of Section B.3.1. As the analysis before, after applying sylow theorem and counting elements, the model reduces the problem to eight possible combinations of Sylow subgroup counts. It then explores a specific case (Combination 1) and derives an intermediate result—that the normalizer of a Sylow 3-subgroup has order 255—but this path does not lead to a contradiction. The model then returns to the general case and successfully deduces a key relationship between the Sylow 17-subgroups and Sylow 3-subgroups. Unfortunately, it then incorrectly applies the intermediate result (order 255), which is valid only for Combination 1, to the general argument, causing the proof to fail.
@@ Appendix F Experiment Details
In Section F.1, we detailed our formal experiment setup, encompassing both the generation and verification phases, while Section F.2 further elaborates on the process, standards, and methodology for our manual natural language evaluations.
@@ F.1.2 Verification
This appendix explained the verification methodology in Section 4.1 employed to ensure the correctness and integrity of the generated Lean proofs, especially considering the models’ flexibility to generate auxiliary definitions and lemmas.
@@
Upon successfully passing all the string-based checks, every formal proof is rigorously checked by the Lean kernel to confirm it contains no sorry or compilation errors. For our experiments, we used different Mathlib versions for each benchmark to align with the state of the library at the time of their creation. Specifically, the evaluations for FATE-M, FATE-H, and FATE-X were conducted using Mathlib versions 4.13.0, 4.16.0, and 4.19.0, respectively. We set a timeout of 300 seconds and use the preloaded Mathlib cache.
@@ F.1.3 Additional Pass@k Results
We report Pass@$k$ results (with $k=2^{n}$ for $n=0,\dots,6$) in Table 9 for the FATE-M and FATE-H benchmarks, omitting FATE-X as no model achieved a successful proof. These results are estimated from $n=64$ samples per problem using the unbiased estimator:
@@
Table 9: Pass@k results across the FATE series.
(a) Pass@k on FATE-M (%).
@@
F.1.4 Performance under Normalized Budget
We report model accuracy using Pass@$k$ under a normalized computational budget. As the input (the question statement) is significantly shorter than the output, we approximate the total cost by model size × average output tokens × number of samples. The results under this normalized budget are shown in Table 10. Closed-source models are excluded due to their undisclosed model sizes, and FATE-X is omitted as all recorded values were zero.
Table 10: Formal accuracy comparison under a normalized sampling budget across the FATE series.
(a) Normalized Accuracy on FATE-M.
Model Size Avg. Length Pass Accuracy (%)
Reasoning model
Deepseek-R1 671B 10700 2 11.3
Qwen3 235B 11700 5 6.9
Theorem Prover
Deepseek-Prover-V2 671B 4600 4 41.0
Goedel-Prover-V2 32B 6300 64 48.7
Kimina-Prover 72B 12000 15 28.0
(b) Normalized Accuracy on FATE-H.
Model Size Avg. Length Pass Accuracy (%)
Reasoning model
Deepseek-R1 671B 17300 1 0.0
Qwen3 235B 17500 3 0.0
Theorem Prover
Deepseek-Prover-V2 671B 11400 2 0.7
Goedel-Prover-V2 32B 13700 30 2.0
Kimina-Prover 72B 21000 9 1.4